what is the answer and way to work this out?
log(base 3)81=x^2

Answer by stanbon(75887)

log(base 3)81=x^2
----
base = 3
log = exponent = x^2
result = 81
----
Equation:
3^(x^2) = 81
---
3^(x^2) = 3^4
---
x^2 = 4
x = 2 or x = -2
--------
Cheers,
Stan H.
==================